OpenStudy (anonymous):

The number of three-digit numbers with distinct digits that can be formed using the digits 1, 2, 3, 5, 8, and 9 is.........? . The probability that both the first digit and the last digit of the three-digit number are even numbers is.........? .

OpenStudy (anonymous):

The number of combinations of \(k\) distinct elements from a set of \(n\) elements is given by\[\left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{ n! }{ k!\left( n-k \right)!}\]In this problem, \(n=6\) and \(k=3\)

OpenStudy (anonymous):

alright thanks...

OpenStudy (anonymous):

nah dude i got the rest... it's all good

OpenStudy (anonymous):

moderator?

OpenStudy (anonymous):

Your welcome. The probability part is straight forward as there are only two possibilities; either the first digit is 2 and the last one is 8, or vice versa

OpenStudy (anonymous):

alright sounds good... you have a goodnight, thanks again @ospreytriple

OpenStudy (anonymous):

moderator... answer me

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

answer me, instead of being creepy watching my post @Kainui